Guile-GNOME: libgnomeui

version 2.16.2, updated 9 December 2011

Kjartan Maraas Anders Carlsson many others

This manual is for (gnome libgnomeui) (version 2.16.2, updated 9 December 2011) Copyright 1998-2007 Kjartan Maraas, Anders Carlsson, and others

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU General Public License, Version 2 or any later version published by the Free Software Foundation.

Short Contents

1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 GnomeAuthentication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 GnomeClient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 4 GnomeDateEdit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5 GnomeMultiScreen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 6 GnomePassword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 7 gnome-vfs-util . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 8 Undocumented . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Type Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Function Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

1 Overview

(gnome gnome-ui) wraps some pieces of the libgnomeui library for Guile. It is a part of Guile-GNOME.

Libgnomeui historically was a staging ground for code meant to go into GTK. In the modern environment, it has largely been replaced, its functionality having been pushed into GTK+ itself.

However, as of 2008 it still has some utility as a kind of glue layer between lower-level modules, with the authentication manager, the GNOME-VFS GDK-Pixbuf integration, and the session manager code.

See the documentation for (gnome gobject) for more information on Guile-GNOME.

2 GnomeAuthentication

2.1 Overview

2.2 Usage

[Function] gnome-authentication-manager-init

This function checks for thread support and does a thread initialisation, if the support is available. Also sets the default sync and async gnome-vfs callbacks for various types of authentication. Note: If you call this, and you use threads with gtk+, you must never hold the gdk lock while doing synchronous gnome-vfs calls. Otherwise an authentication callback presenting a dialog could try to grab the already held gdk lock, causing a deadlock. Since 2.4

3 GnomeClient

Interactions with the session manager.

3.1 Overview

3.2 Usage

[Class] <gnome-client>
Derives from <gtk-object>.

This class defines no direct slots.

[Signal on <gnome-client>] save-yourself (arg0 <gint>)
   (arg1 <gnome-save-style>) (arg2 <gboolean>)
   (arg3 <gnome-interact-style>) (arg4 <gboolean>) ⇒ <gboolean>

Called when either a "SaveYourself" or a "SaveYourselfPhase2" call is made by the session manager.

[Signal on <gnome-client>] die

Called when the session manager wants the client to shut down.

[Signal on <gnome-client>] save-complete

Called when the session manager has finished checkpointing all of the clients. Clients are then free to change their state.

[Signal on <gnome-client>] shutdown-cancelled

Called if the session manager had sent a "SaveYourself" to all clients in preparation for shutting down and the shutdown was then cancelled. A client can then continue running and change its state.

[Signal on <gnome-client>] connect (arg0 <gboolean>)

Called once the client has been connected to the signal manager.

[Signal on <gnome-client>] disconnect

Called when the client is disconnected from the session manager.

[Function] gnome-master-client ⇒ (ret <gnome-client>)

Get the master session management client. This master client gets a client id, that may be specified by the ’–sm-client-id’ command line option. A master client will be generated by gnome-program-init. If possible the master client will contact the session manager after command-line parsing is finished (unless gnome-client-disable-master-connection was called). The master client will also set the SM CLIENT ID property on the client leader window of your application. Additionally, the master client gets some static arguments set automatically (see

gnome-client-add-static-arg for static arguments): gnome-program-init passes

all the command line options which are recognised by gtk as static arguments to the master client.

ret Pointer to the master client

[Function] gnome-client-get-config-prefix (self <gnome-client>)
   ⇒ (ret mchars)
[Method] get-config-prefix

Get the config prefix for a client. This config prefix provides a suitable place to store any details about the state of the client which can not be described using the app’s command line arguments (as set in the restart command). You may push the returned value using gnome-config-push-prefix and read or write any values you require.

client Pointer to GNOME session client object.

ret Config prefix. The returned string belongs to libgnomeui library and should NOT be freed by the caller.

[Function] gnome-client-get-flags (self <gnome-client>)
   ⇒ (ret <gnome-client-flags>)
[Method] get-flags

Determine the client’s status with the session manager.,

client Pointer to GNOME session client object.

ret Various <gnome-client-flag> flags which have been or’d together.

[Function] gnome-client-set-restart-style (self <gnome-client>)
   (style <gnome-restart-style>)
[Method] set-restart-style

Tells the session manager how the client should be restarted in future session. The options are given by the <gnome-restart-style> enum.

client Pointer to GNOME session client object.

style When to restart the client.

[Function] gnome-client-set-priority (self <gnome-client>)
   (priority unsigned-int)
[Method] set-priority

The gnome-session manager restarts clients in order of their priorities in a similar way to the start up ordering in SysV. This function allows the app to suggest a position in this ordering. The value should be between 0 and 99. A default value of 50 is assigned to apps that do not provide a value. The user may assign a different priority.

client Pointer to GNOME session client object.

priority Position of client in session start up ordering.

[Function] gnome-client-set-current-directory (self <gnome-client>)
   (dir mchars)
[Method] set-current-directory

Set the directory to be in when running shutdown, discard, restart, etc. commands.

client Pointer to GNOME session client object.

dir Directory path.

[Function] gnome-client-set-environment (self <gnome-client>)

(name mchars) (value mchars)

[Method] set-environment

Set an environment variable to be placed in the client’s environment prior to running restart, shutdown, discard, etc. commands.

client Pointer to GNOME session client object.

name Name of the environment variable

value Value of the environment variable

[Function] gnome-client-set-process-id (self <gnome-client>) (pid int)
[Method] set-process-id

The client should tell the session manager the result of getpid. However, GNOME does this automatically; so you do not need this function.

client Pointer to GNOME session client object.

pid PID to set as the client’s PID.

[Function] gnome-client-set-program (self <gnome-client>) (program mchars)
[Method] set-program

Used to tell the session manager the name of your program. Set automatically; this function isn’t needed.

client Pointer to GNOME session client object.

program Name of the program.

[Function] gnome-client-set-user-id (self <gnome-client>) (id mchars)
[Method] set-user-id

Tell the session manager the user’s login name. GNOME does this automatically; no need to call the function.

client Pointer to GNOME session client object.

id Username.

[Function] gnome-client-save-any-dialog (self <gnome-client>)
   (dialog <gtk-dialog>)
[Method] save-any-dialog

May be called during a "save youself" handler to request that a (modal) dialog is presented to the user. The session manager decides when the dialog is shown, but it will not be shown it unless the session manager is sending an interaction style of <gnome-interact-any>. A "Cancel Logout" button will be added during a shutdown.

client Pointer to <gnome-client> object.

dialog Pointer to GNOME dialog widget (a <gtk-dialog> widget).

[Function] gnome-client-save-error-dialog (self <gnome-client>)
   (dialog <gtk-dialog>)
[Method] save-error-dialog

May be called during a "save youself" handler when an error has occurred during the save. The session manager decides when the dialog is shown, but it will not be shown it unless the session manager is sending an interaction style of <gnome-interact-any>. A "Cancel Logout" button will be added during a shutdown.

client Pointer to <gnome-client> object.

dialog Pointer to GNOME dialog widget (a <gtk-dialog> widget).

[Function] gnome-client-request-phase-2 (self <gnome-client>)
[Method] request-phase-2

Request the session managaer to emit the "save yourself" signal for a second time after all the clients in the session have ceased interacting with the user and entered an idle state. This might be useful if your app manages other apps and requires that they are in an idle state before saving its final data.

client  A <gnome-client> object.
[Function] gnome-client-request-save (self <gnome-client>)
   (save style <gnome-save-style>) (shutdown bool)
   (interact style <gnome-interact-style>) (fast bool) (global bool)
[Method] request-save

Request the session manager to save the session in some way. The arguments correspond with the arguments passed to the "save yourself" signal handler. The save style indicates whether the save should affect data accessible to other users (<gnome-save-global>) or only the state visible to the current user (<gnome-save-local>) or both. Setting shutdown to ‘#t’ will initiate a logout. The interact style specifies which kinds of interaction will be available. Setting fast to ‘#t’ will limit the save to setting the session manager properties plus any essential data. Setting the value of global to ‘#t’ will request that all the other apps in the session do a save as well. A global save is mandatory when doing a shutdown.

client Pointer to GNOME session client object.

save-style Save style to request.

shutdown Whether to log out of the session.

interact-style Whether to allow user interaction.

fast Minimize activity to save as soon as possible.

global Request that all other apps in the session also save their state.

[Function] gnome-client-flush (self <gnome-client>)

[Method] flush

This will force the underlying connection to the session manager to be flushed. This is useful if you have some pending changes that you want to make sure get committed.

client A <gnome-client> instance.

[Function] gnome-client-new ⇒ (ret <gnome-client>)

Allocates memory for a new GNOME session management client object. After allocating, the client tries to connect to a session manager. You probably want to use

gnome-master-client instead.

ret Pointer to a newly allocated GNOME session management client object.

[Function] gnome-client-new-without-connection ⇒ (ret <gnome-client>)

Allocates memory for a new GNOME session management client object. You probably

want to use gnome-master-client instead.

ret Pointer to a newly allocated GNOME session management client object.

[Function] gnome-client-connect (self <gnome-client>)

[Method] connect

Causes the client to connect to the session manager. Usually happens automatically; no need to call this function.

client A <gnome-client> instance.

[Function] gnome-client-disconnect (self <gnome-client>)
[Method] disconnect

Disconnect the client from the session manager.

client A <gnome-client> instance.

[Function] gnome-client-set-id (self <gnome-client>) (id mchars)

[Method] set-id

Set the client’s session management ID; must be done before connecting to the session manager. There is usually no reason to call this function.

client A <gnome-client> instance.

id Session management ID.

[Function] gnome-client-get-id (self <gnome-client>) ⇒ (ret mchars)

[Method] get-id

Returns session management ID

client A <gnome-client> instance.

ret Session management ID for this client; ‘#f’ if not connected to a session manager.

[Function] gnome-client-get-previous-id (self <gnome-client>)
   ⇒ (ret mchars)
[Method] get-previous-id

Get the session management ID from the previous session.

client A <gnome-client> instance.

ret Pointer to the session management ID the client had in the last session, or ‘#f’ if it was not in a previous session.

[Function] gnome-client-get-desktop-id (self <gnome-client>)
   ⇒ (ret mchars)
[Method] get-desktop-id

Get the client ID of the desktop’s current instance, i.e. if you consider the desktop as a whole as a session managed app, this returns its session ID using a GNOME extension to session management. May return ‘#f’ for apps not running under a recent version of gnome-session; apps should handle that case.

Chapter 3: GnomeClient 8 client A <gnome-client> instance. ret Session ID of GNOME desktop instance, or ‘#f’ if none.

4 GnomeDateEdit

4.1 Overview

4.2 Usage

[Class] <gnome-date-edit>

Derives from <gtk-hbox>. This class defines the following slots:

time The time currently selected

dateedit-flags

Flags for how DateEdit looks

lower-hour

Lower hour in the time popup selector

upper-hour

Upper hour in the time popup selector

initial-time

The initial time

[Signal on <gnome-date-edit>] time-changed
[Signal on <gnome-date-edit>] date-changed

[Function] gnome-date-edit-new (the time long) (show time bool)

(use 24 format bool) ⇒ (ret <gtk-widget>) Creates a new <gnome-date-edit> widget which can be used to provide an easy to use way for entering dates and times. If the-time is 0 then current time is used.

the-time date and time to be displayed on the widget

show-time whether time should be displayed

use-24-format whether 24-hour format is desired for the time display.

ret  a new <gnome-date-edit> widget.
[Function] gnome-date-edit-new-flags (the time long)
   (flags <gnome-date-edit-flags>) ⇒ (ret <gtk-widget>)

Creates a new <gnome-date-edit> widget with the specified flags. If the-time is 0 then current time is used.

the-time The initial time for the date editor.

flags A bitmask of GnomeDateEditFlags values.

ret the newly-created date editor widget.

[Function] gnome-date-edit-construct (self <gnome-date-edit>)
   (the time long) (flags <gnome-date-edit-flags>)
[Method] construct

For language bindings and subclassing only

gde The <gnome-date-edit> object to construct

the-time The initial time for the date editor.

flags A bitmask of GnomeDateEditFlags values.

[Function] gnome-date-edit-set-time (self <gnome-date-edit>)

(the time long)

[Method] set-time

Changes the displayed date and time in the GnomeDateEdit widget to be the one represented by the-time. If the-time is 0 then current time is used.

gde the GnomeDateEdit widget

the-time The time and date that should be set on the widget

[Function] gnome-date-edit-get-time (self <gnome-date-edit>) ⇒ (ret long)
[Method] get-time

gde The GnomeDateEdit widget

ret the time entered in the GnomeDateEdit widget

[Function] gnome-date-edit-set-popup-range (self <gnome-date-edit>)

(low hour int) (up hour int)

[Method] set-popup-range

Sets the range of times that will be provide by the time popup selectors.

gde The GnomeDateEdit widget

low-hour low boundary for the time-range display popup.

up-hour upper boundary for the time-range display popup.

[Function] gnome-date-edit-set-flags (self <gnome-date-edit>)
   (flags <gnome-date-edit-flags>)
[Method] set-flags

Changes the display flags on an existing date editor widget.

gde The date editor widget whose flags should be changed.

flags The new bitmask of GnomeDateEditFlags values.

[Function] gnome-date-edit-get-flags (self <gnome-date-edit>) ⇒ (ret int)
[Method] get-flags

Queries the display flags on a date editor widget.

gde The date editor whose flags should be queried.

ret The current display flags for the specified date editor widget.

[Function] gnome-date-edit-get-initial-time (self <gnome-date-edit>)

⇒ (ret long)

[Method] get-initial-time

Queries the initial time that was set using the <gnome-date-edit-set-time> or during creation

gde The date editor whose initial time should be queried

ret The initial time in seconds (standard time t format)

5 GnomeMultiScreen

5.1 Overview

5.2 Usage

[Function] gnome-url-show-on-screen (url mchars) (screen <gdk-screen>)

⇒ (ret bool) Like gnome-url-show, but ensures that the viewer application appears on screen. url The url to display. Should begin with the protocol to use (e.g. "http:",

"ghelp:", etc)

screen  a <gdk-screen>.

error Used to store any errors that result from trying to display the url. ret ‘#t’ if everything went fine, ‘#f’ otherwise (in which case error will contain

the actual error).

Since 2.6

[Function] gnome-help-display-on-screen (file name mchars) (link id mchars)
   (screen <gdk-screen>) ⇒ (ret bool)

Like gnome-help-display, but ensures that the help viewer application appears on screen. file-name The name of the help document to display. link-id Can be ‘#f’. If set, refers to an anchor or section id within the requested

document.

screen  a <gdk-screen>.

error A <g-error> instance that will hold the specifics of any error which

occurs during processing, or ‘#f’

ret ‘#t’ on success, ‘#f’ otherwise (in which case error will contain the actual

error).

Since 2.6

[Function] gnome-help-display-uri-on-screen (help uri mchars)
   (screen <gdk-screen>) ⇒ (ret bool)

Like gnome-help-display-uri, but ensures that the help viewer application appears on screen. help-uri The URI to display.

screen  a <gdk-screen>.

error A <g-error> instance that will hold the specifics of any error which

occurs during processing, or ‘#f’

ret ‘#t’ on success, ‘#f’ otherwise (in which case error will contain the actual

error).

Since 2.6

6 GnomePassword

6.1 Overview

6.2 Usage

[Class] <gnome-password-dialog>
Derives from <gtk-dialog>.

This class defines the following slots:

show-username

show-username

show-domain

show-domain

show-password

show-password

show-new-password

show-new-password

show-new-password-quality

show-new-password-quality

show-userpass-buttons

show-userpass-buttons

show-remember

show-remember

readonly-username

readonly-username

readonly-domain

readonly-domain

anonymous

anonymous

remember-mode

remember-mode

message message
message-markup

message-markup

username username
domain  domain
password password
new-password

new-password

[Function] gnome-password-dialog-anon-selected
   (self <gnome-password-dialog>) ⇒ (ret bool)
[Method] anon-selected

Checks whether anonymous support is set to #t and the radio button for connecting as anonymous user is active.

password-dialog

          A <gnome-password-dialog>

ret #t if anonymous support is set and the radio button is active, #f otherwise.

[Function] gnome-password-dialog-get-domain
   (self <gnome-password-dialog>) ⇒ (ret mchars)
[Method] get-domain

Gets the domain name from the password dialog.

password-dialog

          A <gnome-password-dialog>

ret The domain name, a char*.

Since 2.4

[Function] gnome-password-dialog-get-password
   (self <gnome-password-dialog>) ⇒ (ret mchars)
[Method] get-password

Gets the password from the password dialog.

password-dialog

          A <gnome-password-dialog>

ret The password, a char*.

Since 2.4

[Function] gnome-password-dialog-get-username
   (self <gnome-password-dialog>) ⇒ (ret mchars)
[Method] get-username

Gets the username from the password dialog.

password-dialog

          A <gnome-password-dialog>

ret The username, a char*.

Since 2.4

[Function] gnome-password-dialog-new (dialog title mchars) (message mchars)

(username mchars) (password mchars) (readonly username bool)

   ⇒ (ret <gtk-widget>)

Creates a new password dialog with an optional title, message, username, password etc. The user will be given the option to save the password for this session only or store it permanently in her keyring. dialog-title

The title of the dialog

message Message text for the dialog

username The username to be used in the dialog

password Password to be used

readonly-username Boolean value that controls whether the user can edit the username or not

ret A new password dialog.

Since 2.4

[Function] gnome-password-dialog-run-and-block
   (self <gnome-password-dialog>) ⇒ (ret bool)
[Method] run-and-block

Gets the user input from PasswordDialog.

password-dialog

          A <gnome-password-dialog>

ret ‘#t’ if "Connect" button is pressed. ‘#f’ if "Cancel" button is pressed.

Since 2.4

[Function] gnome-password-dialog-set-domain
   (self <gnome-password-dialog>) (domain mchars)
[Method] set-domain

Sets the domain field in the password dialog to domain.

password-dialog

          A <gnome-password-dialog>

domain The domain that should be set

Since 2.4

[Function] gnome-password-dialog-set-password
   (self <gnome-password-dialog>) (password mchars)
[Method] set-password

Sets the password in the password dialog.

password-dialog

          A <gnome-password-dialog>

password The password that should be set

Since 2.4

[Function] gnome-password-dialog-set-remember
   (self <gnome-password-dialog>)
   (remember <gnome-password-dialog-remember>)
[Method] set-remember

Based on the value of <gnome-password-dialog-remember>, sets the state of the check buttons to remember password for the session and save password to keyring . password-dialog

          A <gnome-password-dialog>.
remember A <gnome-password-dialog-remember>.

Since 2.6

[Function] gnome-password-dialog-set-username
   (self <gnome-password-dialog>) (username mchars)
[Method] set-username

Sets the username in the password dialog.

password-dialog

          A <gnome-password-dialog>

username The username that should be set

Since 2.4

7 gnome-vfs-util

7.1 Overview

7.2 Usage

[Function] gnome-gdk-pixbuf-new-from-uri (uri mchars)
   ⇒ (ret <gdk-pixbuf>)

Loads a GdkPixbuf from the image file uri points to, scaling it to the desired size. If you pass -1 for width or height then the value specified in the file will be used. When preserving aspect ratio, if both height and width are set the size is picked such that the scaled image fits in a width * height rectangle.

uri the uri of an image

ret The loaded pixbuf, or NULL on error

Since 2.14

8 Undocumented

The following symbols, if any, have not been properly documented.

8.1 (gnome gw libgnomeui)

[Variable] gnome-authentication-manager-dialog-is-visible
[Variable] gnome-client-get-global-config-prefix
[Variable] gnome-client-set-global-config-prefix
[Variable] gnome-gdk-pixbuf-new-from-uri-at-scale
[Function] gnome-help-display-desktop-on-screen
[Function] gnome-help-display-with-doc-id-on-screen
[Variable] gnome-password-dialog-get-remember
[Variable] gnome-password-dialog-set-readonly-domain
[Variable] gnome-password-dialog-set-readonly-username
[Variable] gnome-password-dialog-set-show-domain
[Variable] gnome-password-dialog-set-show-password
[Variable] gnome-password-dialog-set-show-remember
[Variable] gnome-password-dialog-set-show-username
[Variable] gnome-password-dialog-set-show-userpass-buttons

Type Index 19

Type Index

<gnome-client> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 <gnome-date-edit>. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

                                                <gnome-password-dialog>. . . . . . . . . . . . . . . . . . . . . 13

Function Index 20 Function Index 21

Function Index

A

anon-selected . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

C

connect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 connect on <gnome-client> . . . . . . . . . . . . . . . . . . . . 3 construct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

D

date-changed on <gnome-date-edit> . . . . . . . . . . . 9

die on <gnome-client> . . . . . . . . . . . . . . . . . . . . . . . . . 3 disconnect. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

disconnect on <gnome-client> . . . . . . . . . . . . . . . . . 3

F

flush . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

G

get-config-prefix. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 get-desktop-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 get-domain. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 get-flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 10 get-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 get-initial-time. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 get-password . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 get-previous-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 get-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 get-username . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

gnome-authentication-manager-init . . . . . . . . . . 2

gnome-client-connect . . . . . . . . . . . . . . . . . . . . . . . . . 7

gnome-client-disconnect . . . . . . . . . . . . . . . . . . . . . . 7

gnome-client-flush . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

gnome-client-get-config-prefix. . . . . . . . . . . . . . 4
gnome-client-get-desktop-id . . . . . . . . . . . . . . . . . 7

gnome-client-get-flags . . . . . . . . . . . . . . . . . . . . . . . 4 gnome-client-get-id . . . . . . . . . . . . . . . . . . . . . . . . . . 7

gnome-client-get-previous-id . . . . . . . . . . . . . . . . 7

gnome-client-new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

gnome-client-new-without-connection . . . . . . . . 7
gnome-client-request-phase-2 . . . . . . . . . . . . . . . . 6
gnome-client-request-save . . . . . . . . . . . . . . . . . . . 6
gnome-client-save-any-dialog . . . . . . . . . . . . . . . . 5
gnome-client-save-error-dialog. . . . . . . . . . . . . . 5
gnome-client-set-current-directory . . . . . . . . . 4
gnome-client-set-environment . . . . . . . . . . . . . . . . 5

gnome-client-set-id . . . . . . . . . . . . . . . . . . . . . . . . . . 7

gnome-client-set-priority . . . . . . . . . . . . . . . . . . . 4
gnome-client-set-process-id . . . . . . . . . . . . . . . . . 5
gnome-client-set-program. . . . . . . . . . . . . . . . . . . . . 5
gnome-client-set-restart-style. . . . . . . . . . . . . . 4
gnome-client-set-user-id. . . . . . . . . . . . . . . . . . . . . 5
gnome-date-edit-construct . . . . . . . . . . . . . . . . . . 10
gnome-date-edit-get-flags . . . . . . . . . . . . . . . . . . 10

gnome-date-edit-get-initial-time . . . . . . . . . . 11
gnome-date-edit-get-time . . . . . . . . . . . . . . . . . . . 10

gnome-date-edit-new . . . . . . . . . . . . . . . . . . . . . . . . . . 9

gnome-date-edit-new-flags . . . . . . . . . . . . . . . . . . . 9
gnome-date-edit-set-flags . . . . . . . . . . . . . . . . . . 10
gnome-date-edit-set-popup-range . . . . . . . . . . . 10
gnome-date-edit-set-time . . . . . . . . . . . . . . . . . . . 10
gnome-gdk-pixbuf-new-from-uri . . . . . . . . . . . . . . 17
gnome-help-display-desktop-on-screen. . . . . . 18
gnome-help-display-on-screen . . . . . . . . . . . . . . . 12
gnome-help-display-uri-on-screen . . . . . . . . . . 12
gnome-help-display-with-doc-id-on-screen

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 gnome-master-client . . . . . . . . . . . . . . . . . . . . . . . . . . 3

gnome-password-dialog-anon-selected . . . . . . . 14
gnome-password-dialog-get-domain . . . . . . . . . . 14
gnome-password-dialog-get-password . . . . . . . . 14
gnome-password-dialog-get-username . . . . . . . . 14
gnome-password-dialog-new . . . . . . . . . . . . . . . . . . 14
gnome-password-dialog-run-and-block . . . . . . . 15
gnome-password-dialog-set-domain . . . . . . . . . . 15
gnome-password-dialog-set-password . . . . . . . . 15
gnome-password-dialog-set-remember . . . . . . . . 15
gnome-password-dialog-set-username . . . . . . . . 16
gnome-url-show-on-screen . . . . . . . . . . . . . . . . . . . 12

R

request-phase-2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 request-save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 run-and-block . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

S

save-any-dialog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

save-complete on <gnome-client> . . . . . . . . . . . . . 3

save-error-dialog. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

save-yourself on <gnome-client> . . . . . . . . . . . . . 3

set-current-directory . . . . . . . . . . . . . . . . . . . . . . . . 4 set-domain. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 set-environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 set-flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 set-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 set-password . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 set-popup-range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 set-priority . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 set-process-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 set-program. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 set-remember . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 set-restart-style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 set-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 set-user-id. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 set-username . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

shutdown-cancelled on <gnome-client>. . . . . . . . 3

T

time-changed on <gnome-date-edit> . . . . . . . . . . . 9